# iPanda: An Intelligent Protocol Testing and Debugging Agent for Conformance Testing

### v1

@@ Abstract.
Conformance testing is essential for ensuring that protocol implementations comply with their specifications. However, traditional testing approaches involve manually creating numerous test cases and scripts, making the process labor-intensive and inefficient. Recently, Large Language Models (LLMs) have demonstrated impressive text comprehension and code generation abilities, providing promising opportunities for automation. In this paper, we propose iPanda, the first end-to-end framework that leverages LLMs to automate protocol conformance testing. Given a protocol specification document and its implementation, iPanda first employs a keyword-based method to automatically generate comprehensive test cases. Then, it utilizes a code-based retrieval-augmented generation approach to effectively interpret the implementation and produce executable test code. To further enhance code quality, iPanda incorporates an iterative self-correction mechanism to refine generated test scripts interactively. Finally, by executing and analyzing the generated tests, iPanda systematically verifies compliance between implementations and protocol specifications. Comprehensive experiments on various protocols show that iPanda significantly outperforms pure LLM-based approaches, improving the success rate () of test-code generation by factors ranging from to .
1. Introduction
@@
Communication protocols such as HTTP, MQTT, and CoAP form the backbone of today’s information-driven society, playing critical roles in internet data transmission, IoT connectivity, and cloud computing services. Ensuring the stability and efficiency of protocol implementations is therefore paramount and necessitates rigorous testing procedures. Conformance testing, specifically designed to verify adherence of protocol implementations to their official standards and specifications, is one of the essential methods employed in validating communication protocols. In typical scenarios, testers might spend weeks manually writing extensive test scripts utilizing protocol implementation libraries, and individually verifying compliance with protocol requirements. For instance, a typical compliance check of MQTT protocol implementations might require engineers to manually develop hundreds of script-based tests, each verifying different message formats and states. With the increasing complexity of protocols, traditional manual testing methods have become inefficient, cumbersome, and difficult to generalize. This situation underscores an urgent need for automated testing solutions that minimize human intervention while enhancing efficiency and coverage.
@@
Concurrently, large language models (LLMs) have shown impressive capabilities in language comprehension, reasoning, and performing sophisticated tasks. Trained on extensive datasets, LLMs effectively generalize to novel tasks, parse complex inputs, and maintain context, making them ideal for automating intricate operations. For instance, in IoT fuzz testing, LLM-driven automation significantly enhances protocol message generation, increasing vulnerability detection effectiveness and uncovering previously undetected issues (Wang et al., 2024). Similarly, in mobile device automation, LLMs combine general reasoning with domain-specific expertise, facilitating complex tasks without extensive manual scripting (Wen et al., 2024a, b). These examples demonstrate the transformative potential of LLMs in turning labor-intensive procedures into efficient automated solutions.
@@
Given this context, a compelling question arises: Can we leverage the capabilities of LLMs to address the challenges inherent in conformance testing of communication protocols? However, the practical realization of this vision faces several significant hurdles:
Efficiently generating comprehensive test cases that rigorously satisfy protocol testing requirements remains a significant and challenging task.
Equipping LLMs with the capability to accurately interpret, adapt, and interact with existing protocol implementation libraries poses considerable difficulty, due to varying specifications and complexity.
Improving the overall performance of LLM-driven testing processes is vital, especially regarding the accuracy, effectiveness, and reliability of the generated test code, ensuring conformance and dependability in real-world testing.
@@
Motivated by trust in the adaptive reasoning capabilities of LLMs and inspired by successful applications in related domains, we tackle the aforementioned challenges by introducing iPanda, an Intelligent Protocol Testing and Debugging Agent. To the best of our knowledge, iPanda represents the first end-to-end, LLM-based solution specifically designed for automating communication protocol conformance testing. It streamlines the entire testing process—from protocol specification to result analysis—significantly reducing manual effort and improving overall efficiency.
Specifically, iPanda leverages large language models to autonomously generate comprehensive test cases directly from protocol specification documents. Given a specific protocol implementation library, iPanda dynamically generates executable test-case code, runs these tests within a simulated environment, and systematically identifies potential compliance issues. To efficiently generate test cases aligned with precise testing requirements, we propose a novel keyword-based test-case generation approach grounded in the inherent characteristics of protocol specifications.
Moreover, to ensure the rapid adaptation of iPanda’s core (i.e., the underlying LLMs) to varying protocol libraries, we developed a code-based retrieval-augmented generation (RAG) mechanism. This mechanism guides the agent dynamically and accurately in employing diverse protocol implementation libraries. Additionally, to further enhance the accuracy and reliability of the generated test code, we incorporate a code self-correction mechanism. This allows the LLM to iteratively validate and refine its outputs interactively, closely mimicking human debugging procedures.
iPanda also supports natural-language commands and demonstrates reasoning capabilities for complex state transitions, making it particularly effective for dynamic, heterogeneous network environments where traditional testing methods frequently fall short. By significantly reducing the dependency on human expertise and fully automating conformance test execution, iPanda presents a transformative advancement in communication protocol conformance testing.
@@
To the best of our knowledge, we present the first automated protocol conformance testing framework integrating LLMs with domain-specific expertise, including keyword-based test case generation, code-based RAG, and a self-correction mechanism for test-code refinement.
@@
Comprehensive experiments demonstrate that, compared to the pure LLM-based method, iPanda improves the of test code generation by to .
Upon acceptance of this paper, we will publicly release the associated source code and datasets to facilitate future research and practical applications.
2. Background and motivation
2.1. Conformance testing of protocols
@@
Test case implementation. Write test scripts or code using the protocol implementation under test to concretely realize the test cases.
Test execution. Run the test case code in a properly configured testing environment and collect results.
@@
2.2. Large language models
Large Language Models (LLMs) are artificial intelligence models built on deep learning techniques, pre-trained on vast amounts of textual data to capture rich semantic and contextual relationships in language. Prominent LLMs, such as the GPT series (Jaech et al., 2024; Achiam et al., 2023; Hurst et al., 2024), LLaMA series (Touvron et al., 2023; Grattafiori et al., 2024), and DeepSeek series (Guo et al., 2025; Liu et al., 2024), possess the capability to generate detailed content, understand complex semantics, and perform a wide range of tasks, including text generation, question answering, and code writing.
To better leverage LLM capabilities, researchers introduced prompt engineering, carefully designing prompts or instructions to guide models toward generating more accurate and effective outputs. Building upon this, a series of prompt augmentation techniques have emerged, further improving model adaptability and generalization. This work specifically adopts several such methods: Chain of Thought (CoT) (Wei et al., 2022), instructing models to explicitly provide intermediate reasoning steps to better handle complex reasoning tasks; In-Context Learning (ICL) (Dong et al., 2022), guiding models with example-based formats to rapidly adapt without additional training; Retrieval-Augmented Generation (RAG) (Lewis et al., 2020), integrating external knowledge or retrieval mechanisms to enhance performance on knowledge-intensive tasks; and Self-Correcting with Tool-Interactive Critiquing (CRITIC) (Alinezhad et al., 2019), iteratively refining outputs through external tool-driven feedback. Together, these approaches significantly boost LLM effectiveness and adaptability in real-world applications.
2.3. Integrating LLM and conformance testing
@@
The current conformance testing process has significant limitations, primarily due to its heavy reliance on manual effort. Conformance testing involves a vast number of detailed test cases that must be manually written, making the process labor-intensive and time-consuming. Additionally, implementing these test cases requires extensive test code development, further increasing the cost of development and maintenance. These limitations significantly reduce the efficiency and flexibility of conformance testing.
@@
At present, no end-to-end conformance testing tool exists that seamlessly integrates protocol documents, implementations, and test results. The exceptional text comprehension and code generation capabilities of LLMs offer new possibilities for addressing these shortcomings. By leveraging the powerful language understanding and generation abilities of LLMs, test case design and test code development can be automated or semi-automated, substantially reducing the manual workload. Therefore, this work aims to integrate LLMs into the conformance testing process for communication protocols. We propose an end-to-end conformance testing agent framework to enhance the efficiency and effectiveness of conformance testing.
3. Design of iPanda
@@
In this section, we provide a detailed introduction to the design of iPanda. This system is specifically designed for communication protocol conformance testing, particularly in network environments such as the Internet of Things. iPanda can dynamically generate test case sets for conformance testing based on protocol specification documents. For a given protocol implementation library, it can generate test case code, interact with the implementation in a simulated communication environment, perform testing and debugging, and analyze execution results to identify potential deficiencies in the implementation.
3.1. Overview
@@
The overview of iPanda is shown in Fig. 1. Taking the RSocket protocol (rsocket, 2024) as an example, suppose a user wants to verify whether its Python implementation rsocket-py (rsocket py, 2025) conforms to its specification. iPanda first extracts key functional points from the protocol document and automatically generates standardized test cases using the LLM generator, optionally applying a filter to remove anomalous cases. For each test case, iPanda generates executable test code using the target implementation library. It retrieves relevant context from the implementation library, integrating this context into detailed prompts that clearly define the LLM’s role and task objectives. To ensure executability, the generated code undergoes validation; if issues are detected, iPanda initiates iterative refinement using historical context and error information, dynamically adjusting prompts to debug the code. Once the generation yields executable code or reaches a retry limit, iPanda compiles a final debugging report and evaluates test outcomes, determining compliance with the protocol’s conformance requirements.
@@
To address the challenges mentioned in Sec. 1 and enhance iPanda’s effectiveness in test cases and code generation, we have designed several optimization methods within iPanda. These techniques are embedded within individual modules or integrated across multiple modules. The following sub-sections will provide a detailed explanation of these optimizations.
3.2. Test case generation
3.2.1. Test case generation methods
Traditional test case generation methods include:
Specification-driven methods. Based on the protocol’s standard specification, these methods analyze state machines and message interaction flows to design test cases that cover both valid and invalid scenarios.
Model-driven methods. By constructing formal models such as finite state machines and extended finite state machines, these methods automatically generate test sequences to ensure comprehensive coverage of states and message transitions.
Implementation-driven methods. These methods test the completeness and robustness of a protocol implementation by analyzing code coverage and injecting faults based on the specific implementation.
Interoperability testing methods. These methods aim at verifying compatibility between different vendors and versions of a protocol, ensuring correct communication between systems.
Random and fuzz testing methods. These methods use random messages and anomalous data inputs to uncover hidden defects in the protocol implementation, improving security and fault tolerance.
The above methods often require significant human and material resources from professionals, lacking sufficient flexibility and scalability. For example, in the model-driven methods, constructing a formal model specific to a communication protocol is necessary, but this model cannot be easily extended to other protocol conformance testing tasks.
Given LLMs’ great text comprehension and generation capabilities, they are naturally suited for analyzing and extracting information from specification documents. For instance, even without fine-tuning, an LLM can identify key statements from a given RFC document simply by being prompted to extract relevant textual information. However, due to the inherent randomness of LLMs, they do not strictly adhere to instructions, making extracted information potentially nonconformant or uncontrolled. Therefore, solely relying on LLMs for the direct extraction of key information is impractical and unreliable.
3.2.2. Keyword-based test case generation method
@@
Taking CoAP as an example, we reviewed a total of RFC documents related to CoAP and found that of them comply with RFC 2119 (Shelby et al., 2014; Editor, 2025). These documents use uppercase keywords such as MUST, MUST NOT, REQUIRED, SHALL, and SHALL NOT, to define and emphasize critical protocol specifications and requirements. Even those documents that do not strictly follow RFC 2119 still use similar auxiliary verbs to highlight protocol constraints.
@@
Inspired by this insight, we design a novel specification-driven test case generation method named keyword-based test case generation (keyword-based TCG). This method is integrated into the test case generation module, as shown in Fig. 1. Keyword-based TCG combines heuristic rules with the idea of generating datasets using LLMs. Specifically, this method first detects whether the document follows RFC 2119. If it does, keyword-based TCG uses regular expressions to extract paragraphs containing uppercase keywords. Each extracted paragraph is a complete natural paragraph to preserve as much contextual semantic information as possible. We define these paragraphs as functional points. If the document does not conform to RFC 2119, the method defaults to case-insensitive keyword extraction and repeats the above process.
@@
After extracting functional points, keyword-based TCG calls the LLM to generate test cases based on these points. To ensure a standardized format for the generated test cases, we introduce few-shot in-context learning (few-shot ICL) into the LLM. Specifically, we construct a prompt with input-output example , where represents a functional point, and is a test case in the desired format. These input-output pairs are concatenated in the format:. During reasoning, the test is appended to the prompt, and the LLM learns the structure from the provided examples, generating an output in the same format, as illustrated in Fig. 2. To ensure the generated test cases are both domain-relevant and effective, we introduce a filter, allowing users to review the functional points and their corresponding generated test cases, removing any anomalous test cases.
@@
Remark: It is important to note that the test case generation module is optional, meaning that test cases do not necessarily have to be generated by this module. iPanda allows users to import their own test cases, meanwhile still providing compatibility with subsequent code generation and conformance verification functionalities. The main purposes of the test case generation module are twofold: to facilitate the rapid generation of effective test cases directly from protocol documents, reducing the workload for users; and to provide standardized experimental datasets for this work, enabling a quantitative evaluation of iPanda’s performance.
3.3. Code-oriented prompting
3.3.1. Potential issues in LLMs

### v2

# iPanda: An LLM-based Agent for Automated Conformance Testing of Communication Protocols

@@ Abstract
Conformance testing is essential for ensuring that protocol implementations comply with their specifications. However, traditional testing approaches involve manually creating numerous test cases and scripts, making the process labor-intensive and inefficient. Recently, Large Language Models (LLMs) have demonstrated impressive text comprehension and code generation abilities, providing promising opportunities for automation. In this paper, we propose iPanda, the first framework that leverages LLMs to automate protocol conformance testing. Given a protocol specification document and its implementation, iPanda first employs a keyword-based method to automatically generate comprehensive test cases. Then, it utilizes retrieval-augmented generation and customized CoT strategy to effectively interpret the implementation and produce executable test programs. To further enhance programs’ quality, iPanda incorporates an iterative optimization mechanism to refine generated test scripts interactively. Finally, by executing and analyzing the generated tests, iPanda systematically verifies compliance between implementations and protocol specifications. Comprehensive experiments on various protocols show that iPanda significantly outperforms pure LLM-based approaches, improving the success rate () of test-program generation by factors ranging from to .
Index Terms:
Conformance Testing, LLMs, Automated Testing.I Introduction
@@
Communication protocols such as MQTT, CoAP, WebSocket form the backbone of today’s information-driven society, playing critical roles in internet data transmission, IoT connectivity, and cloud computing services. Ensuring the stability and efficiency of protocol implementations is therefore paramount and necessitates rigorous testing procedures. Conformance testing, specifically designed to verify adherence of protocol implementations to their official standards and specifications, is one of the essential methods employed in validating communication protocols. In typical scenarios, testers might spend weeks manually writing extensive test scripts utilizing protocol implementation libraries, and individually verifying compliance with protocol requirements. For instance, a typical compliance check for the CoAP protocol involves developing hundreds of script-based tests, to verify its message grammar, stateful interactions and operational sequences, as mandated by over MUST and SHOULD requirements in RFC7252. With the increasing complexity of protocols, traditional manual testing methods have become inefficient, cumbersome, and difficult to generalize. This situation underscores an urgent need for automated testing solutions that minimize human intervention while enhancing efficiency and coverage.
@@
Concurrently, large language models (LLMs) have shown impressive capabilities in language comprehension, reasoning, and performing sophisticated tasks. Trained on extensive datasets, LLMs effectively generalize to novel tasks, parse complex inputs, and maintain context, making them ideal for automating intricate operations. For instance, in IoT fuzz testing, LLM-driven automation enhances protocol message generation, increasing vulnerability detection effectiveness and uncovering previously undetected issues [1]. Similarly, in mobile device automation, LLMs combine general reasoning with domain-specific expertise, facilitating complex tasks without extensive manual scripting [2, 3]. These examples demonstrate the transformative potential of LLMs in turning labor-intensive procedures into efficient automated solutions.
@@
Given this context, a compelling question arises: Can we leverage the capabilities of LLMs to address the challenges inherent in conformance testing of communication protocols? Realizing this vision in practice, however, entails overcoming several key obstacles:
Specification-to-test-case derivation. In contrast to fuzzing, conformance testing is intrinsically driven by specifications, requiring the translation of often ambiguous textual specifications into structured, executable test logic.
Knowledge adaptation and grounding. LLMs struggle to bridge the gap between abstract rules described in test cases and the concrete APIs provided by specific implementation libraries, making it difficult to generate valid and context-aware test programs.
Practical executability and reliability. Beyond syntactic correctness, test programs must be semantically aligned with the target library’s API to ensure they are both executable and trustworthy in real-world testing scenarios.
@@
While LLMs have achieved remarkable success in domains such as code generation and dialogue systems, they fall short in addressing the unique demands of protocol conformance testing. We bridge this gap with iPanda, an Intelligent Protocol Testing and Debugging Agent. As the first LLM-based solution for automating communication protocol conformance testing, iPanda streamlines the entire workflow. It autonomously parses a protocol specification to generate comprehensive test cases, translates them into executable programs for a target implementation library, and then executes these tests to systematically identify compliance violations, thereby significantly reducing manual effort.
Especially, we introduce multiple mechanisms in iPanda to overcome the above challenges. To efficiently generate test cases aligned with precise testing requirements, we propose a novel keyword-based test-case generation approach grounded in the inherent characteristics of protocol specifications. It automatically locates the specific specifications that need to be tested based on keywords, and generates formatted test cases. Moreover, to ensure the adaptation of iPanda to varying protocol libraries, we introduce code-oriented retrieval-augmented generation (RAG), which is deeply coupled with the test program generation process. This mechanism iteratively queries the knowledge base at different stages of program synthesis, providing just-in-time, context-specific guidance. Additionally, to enable LLMs to handle complex test cases, we design a customized chain-of-thought (CoT) strategy that operationalizes a divide-and-conquer methodology. It guides iPanda to first decompose the complex task case into logical sub-tasks and solve them individually, before integrating these modular solutions into a complete, high-quality test program. Finally, We introduce testing platform and iterative optimization mechanism to achieve automated testing and program correction, ensuring the program’s executability and reliability.
iPanda also supports natural-language commands and demonstrates reasoning capabilities for complex state transitions, making it effective for dynamic, heterogeneous network environments where traditional methods frequently fall short. By significantly reducing the dependency on human expertise and fully automating test execution, iPanda presents a transformative advancement in protocol conformance testing.
@@
To the best of our knowledge, we present the first automated protocol conformance testing framework integrating LLMs with domain-specific expertise, including keyword-based test case generation, customized CoT strategy, and iterative optimization mechanism for test-program refinement.
@@
As an example, we using iPanda to perform conformance testing on CoAP and RSocket, and present two test case sets, i.e., CoAP-set and RSocket-set. Comprehensive experiments demonstrate that, compared to the pure LLM-based method, iPanda improves the of test program generation by to .
Upon acceptance of this paper, we will publicly release the associated source code and datasets.
II Background and motivation
II-A Conformance Testing of Protocols
@@
Test case implementation. Write test scripts or programs using the protocol implementation under test to concretely realize the test cases.
Test execution. Run the test case programs in a properly configured testing environment and collect results.
@@
II-B Uniqueness of Conformance Testing
Traditional testing, like unit testing and integration testing, is fundamentally implementation-driven. Whether verifying individual code components (unit testing) or the interactions between modules (integration testing), the design, execution, and evaluation of test cases are all rooted in the implementation’s source code and architecture. In short, it aims to ensure that the existing implementation works as expected. In contrast, specification-based conformance testing differs fundamentally in its methodology. Its goal is to verify whether an implementation strictly adheres to an external, authoritative specification (e.g., a technical standard or an official API document). Its uniqueness is reflected mainly in two aspects:
Shift in the source of truth. In conformance testing, the source of truth shifts from the internal code to the external specification. It does not concern itself with implementation details but focuses solely on whether the behavior aligns with what is described in the specification.
Systematic specification-driven process. Conformance testing requires a top-down process. It begins with systematically interpreting and decomposing the specification text into verifiable requirements. Based on these extracted requirements, test cases are then generated and executed to map precisely to actual API calls.
Essentially, traditional testing verifies what the code does, whereas conformance testing verifies whether the code does what the specification requires it to do. This makes conformance testing a systematic endeavor that demands holistic understanding (inferring the functional points to be tested from the specification), careful planning (designing tests that can validate these points), and rigorous execution (generating test programs to implement these tests). It is not merely a straightforward check of code functionality. For this reason, traditional testing tools are inadequate for conformance testing.
II-C Integrating LLM and Conformance Testing
@@
The current conformance testing process has significant limitations, primarily due to its heavy reliance on manual effort. It involves a vast number of detailed test cases that must be manually written, making the process labor-intensive and time-consuming. Additionally, implementing these test cases requires extensive test program development, further increasing the cost of development and maintenance. These limitations significantly reduce the efficiency and flexibility of conformance testing.
@@
At present, no conformance testing tool exists that seamlessly integrates protocol documents, implementations, and test results. The exceptional text comprehension and code generation capabilities of LLMs offer new possibilities for addressing these shortcomings. By leveraging the powerful abilities of LLMs, test case design and test program development can be automated or semi-automated, substantially reducing the manual workload. Therefore, this work aims to integrate LLMs into the conformance testing process for communication protocols. We propose an LLM-based conformance testing agent framework to enhance the efficiency and effectiveness of conformance testing.
III Design of iPanda
@@
In this section, we present the design of iPanda in detail, which is specifically designed for conformance testing of communication protocol. iPanda can dynamically generate test case sets for conformance testing based on protocol specification documents. For a given protocol implementation library, it can generate test case programs, interact with the implementation in a simulated communication environment, perform testing and debugging, and analyze execution results to identify potential deficiencies in the implementation.
@@
The overview of iPanda is shown in Fig. 1. Taking the RSocket protocol [4] as an example, suppose a user wants to verify whether its Python implementation rsocket-py [5] conforms to its specification. iPanda first extracts key functional points from the protocol document and automatically generates standardized test cases using the LLM generator, optionally applying a filter to remove anomalous cases. For each test case, guiding by customized CoT strategy, iPanda generates executable test program using the target implementation library. To generate high-quality program, it retrieves relevant context from the implementation library. To ensure executability, the generated program undergoes validation; if issues are detected, iPanda initiates iterative refinement using historical context and error information, dynamically adjusting prompts to debug the program. Once the generation yields executable program or reaches a retry limit, iPanda compiles a final debugging report and evaluates test outcomes, determining compliance with the protocol’s conformance requirements.
@@
To address the challenges mentioned in Sec. I and enhance iPanda’s effectiveness, we have designed several optimization methods. The following sub-sections will provide detailed descriptions of these methods.
III-A Keyword-based Test Case Generation
In conformance testing, the generation of test cases has traditionally relied heavily on manual effort. For example, in the FIDO2 Conformance Test Tool for the FIDO protocol, developers manually extract as many test requirements as possible from the specification documents and then write corresponding test cases to integrate into the tool[6]. Although some efforts have emerged to automate the parsing of specifications (For instance, RFCSCAN attempts to treat individual RFC sections as test units and directly compare them with the implementation[7]), these approaches remain limited to coarse-grained, section-level parsing. As a result, they lack the capability for more fine-grained analysis of the documents, making it challenging to automatically generate test cases targeting specific functional points.
@@ III-A Keyword-based Test Case Generation
Taking CoAP as an example, we reviewed a total of RFC documents related to CoAP and found that of them comply with RFC 2119 [8, 9]. These documents use uppercase keywords such as MUST, REQUIRED, SHALL, to define and emphasize critical protocol specifications and requirements (named the functional points). Even the documents that do not strictly follow RFC 2119 still use similar auxiliary verbs as keywords to highlight protocol functional points.
Preliminary Experiment: Furthermore, we examine the coverage when using keywords to locate functional points. Assuming the keyword set is , and following RFCSCAN, we take the smallest subsections as the evaluation units. The section coverage is then defined as follows:
where the function measures the string length of , since sections vary in length and longer sections are more likely to contain additional functional points. is the indicator function, defined as follows:
The experiment shows that for 32 CoAP-related RFC documents (considering only the main body), the overall section coverage reaches , with 10 documents achieving a section coverage of . These results demonstrate that using keywords to locate functional points can parse documents at a finer granularity while ensuring high section coverage.
@@
Inspired by the above insight and experiment, we design a novel specification-driven test case generation method named keyword-based test case generation (keyword-based TCG). This method is integrated into the test case generation module, as shown in Fig. 1. Keyword-based TCG combines heuristic rules with the idea of generating datasets using LLMs. Specifically, this method first detects whether the document follows RFC 2119. If it does, keyword-based TCG uses regular expressions to extract paragraphs containing uppercase keywords; if not, it defaults to case-insensitive keyword extraction. Each extracted paragraph is a complete natural paragraph to preserve as much contextual semantic information as possible. We define these paragraphs as functional points.
@@
Based on these functional points, keyword-based TCG calls the LLM to generate test cases. To ensure a standardized format for the test cases, we introduce few-shot in-context learning into the LLM. Specifically, we construct a prompt with input-output example , where represents a functional point, and is a standardized test case. These input-output pairs are concatenated in the format:. During reasoning, the test is appended to the prompt, and the LLM learns the structure from the provided examples, generating an output in the same format, as illustrated in Fig. 2. The test cases generated by the LLM comprise four components: test preconditions, test steps, test assertions, and precautions. The test steps describe in detail the procedure for testing the selected functional point, and the test assertions define the criteria for judging test success, which serve as the basis for evaluating the results. To ensure the test cases are both domain-relevant and effective, we adopt standard processes of data generation [10], introducing a filter to review the functional points and their corresponding test cases, removing any anomalous test cases.
@@
Remark: It is important to note that the test case generation module is optional, meaning that test cases do not necessarily have to be generated by this module. iPanda allows users to import their own test cases, meanwhile still providing compatibility with subsequent program generation and conformance verification functionalities. The main purposes of the test case generation module are twofold: to facilitate the rapid generation of effective test cases directly from protocol documents, reducing the workload for users; and to provide standardized experimental datasets for this work, enabling a quantitative evaluation of iPanda’s performance.
III-B Automated Program Synthesis